OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

y=3x-14

OpenStudy (anonymous):

i can explain if you need me to

OpenStudy (anonymous):

@babygirl11271963 thank you that'd help a lot

OpenStudy (anonymous):

ok well you have to put it in the proper form. 3x-y=-3. move 3x to the other side. -y=-3x-3. isolate y . y=3x+3 (you should have just divided the entire equation by -1 to make y possitive) to be parallel it should have the same slope so you keep y=3x. pluh in your x and y values. -2=3(4)+? -2=12+? -14 so you get y=3x-14

OpenStudy (anonymous):

put in ***

OpenStudy (anonymous):

to be parallel it must have same slope different y intercept. if they were to both be the same then they would be the same line. if you changed the slope and not the intercept, then the y int is where they would intercept.

OpenStudy (anonymous):

Oh okay thank you!